# The Flea on the Elephant: tidal Love numbers in subsolar primordial black hole searches

### v1

@@ I Introduction
The detection of a subsolar mass object in a black hole (BH) merger is regarded as one of the definitive proof of the primordial origin of a binary Green and Kavanagh (2021). However, leveraging this opportunity requires the ability to differentiate such an event from other astrophysical systems and possible new physics candidates Cardoso and Pani (2019); Barsanti et al. (2022); Franciolini et al. (2022); Crescimbeni et al. (2024).
@@
A key distinction between material compact objects and BHs is that the tidal deformability – measured by tidal Love numbers (TLNs) – of the latter is (supposed to be) zero in four-dimensional space-times Deruelle and Piran (1984); Binnington and Poisson (2009); Damour and Nagar (2009); Damour and Lecian (2009); Pani et al. (2015a, b); Gürlebeck (2015); Porto (2016); Le Tiec and Casals (2021); Chia (2021); Le Tiec et al. (2021); Hui et al. (2021); Charalambous et al. (2021a, b); Creci et al. (2021); Bonelli et al. (2022); Ivanov and Zhou (2023a); Charalambous et al. (2022); Katagiri et al. (2023); Ivanov and Zhou (2023b); Berens et al. (2023); Bhatt et al. (2023); Sharma et al. (2024); Rai and Santoni (2024). This unique property of BHs in general relativity can be explained by special symmetries of the equation of motions for static and linear perturbations Hui et al. (2021); Charalambous et al. (2021a, b); Hui et al. (2022a, b); Charalambous et al. (2022); Ivanov and Zhou (2023b); Katagiri et al. (2023); Bonelli et al. (2022); Kehagias et al. (2023); Ben Achour et al. (2022); Berens et al. (2023); Rai and Santoni (2024); Kehagias et al. (2024); Sharma et al. (2024) (see also Refs. De Luca et al. (2023a); Riva et al. (2024); Ivanov et al. (2024); Nair et al. (2023); Saketh et al. (2024); Perry and Rodriguez (2023); Chakraborty et al. (2023); De Luca et al. (2024) for recent studies discussing the role of nonlinearities and time-dependent tidal fields), but it does not apply to any other material object Porto (2016); Cardoso et al. (2017).
@@
This property is however invalidated in scenarios involving the presence of a cosmological constant Nair et al. (2024), in theories of modified gravity Cardoso et al. (2017, 2018); De Luca et al. (2023b); Barura et al. (2024) or higher dimensions Kol and Smolkin (2012); Cardoso et al. (2019); Hui et al. (2021); Rodriguez et al. (2023); Charalambous and Ivanov (2023); Charalambous (2024a, b); Ma et al. (2024), and especially in the presence of an external environment. The latter significantly alters the nature of the tidal interactions, potentially leading to nonzero TLNs for non-vacuum binary BHs. Examples of this effect include BHs surrounded by clouds of ultralight bosonic fields generated through accretion or superradiance Baumann et al. (2019); De Luca and Pani (2021); De Luca et al. (2023c); Brito and Shah (2023), or by fluids of matter and accretion disks Cardoso and Duque (2020); Cardoso et al. (2022); Cannizzaro et al. (2024).
@@
Since TLNs affect the gravitational wave (GW) signal of binary systems at high post-Newtonian (PN) order Flanagan and Hinderer (2008); Hinderer et al. (2010) and could be measured with good accuracy in the presence of large tidal interactions Chia et al. (2023, 2024), they provide a natural key feature to test the nature of the compact object.
@@
Suppose now that in the future a subsolar merger event is detected with a large TLN. One would be immediately driven to think that such an event does not involve a primordial black hole (PBH). We show in this paper that this is not the case. Indeed, even a small environmental disturbance (the “flea”) to a BH (the “elephant”) generates a large and measurable effective TLN of the system. Such a “flea on the elephant” effect, here named in analogy with the quantum mechanical effect Graffi et al. (1984), is not surprising, it appears as well in the spectrum of the quasinormal modes of the GWs emitted during the ringdown phase following the merger of two BHs when tiny perturbations are present Jaramillo et al. (2021); Destounis et al. (2021); Gasperin and Jaramillo (2022); Boyanov et al. (2023); Jaramillo (2022); Kyutoku et al. (2023); Sarkar et al. (2023); Destounis et al. (2024); Areán et al. (2023); Cownden et al. (2024); Destounis and Duque (2023); Courty et al. (2023); Boyanov et al. (2024); Cao et al. (2024); Cardoso et al. (2024); Ianniccari et al. (2024).
@@ II An Effective tidal Love number
The Dirac delta potential $V_{1}$, describing the external bump, is located at a distance $L$ much larger than $r_{s}$, and is the same considered in the aforementioned literature regarding the quasinormal modes. It is supposed to mimic the presence of a matter shell whose density amplitude is ${\cal O}(\epsilon)$, with width ${\cal O}(r_{s})\ll L$ and corresponding mass $\sim\epsilon\left(L/r_{s}\right)m$, where $m$ is the BH mass Cheung et al. (2022). The source creating the tidal force is at distances $\gg L$, see Fig. 1 for a pictorial illustration. This simple toy model captures possible environmental effects present around the PBH system Baumann et al. (2019); De Luca and Pani (2021); De Luca et al. (2023c); Brito and Shah (2023); Cardoso and Duque (2020); Cardoso et al. (2022); Cannizzaro et al. (2024), mimicking for example a layer of a more complicated matter distribution which might thus spread even over larger distances. Our results are therefore intended to be conservative as far as the TLN is concerned.
@@
Thus, the TLNs of PBHs surrounded by weak rings could be sufficiently large and potentially impact the evolution of coalescing binaries. To get the feeling of the numbers, for the characteristic values $\epsilon\sim 10^{-4}$ and $L\sim 30\,r_{s}$, one obtains sizeable TLNs as $k_{2}\sim-1.5\cdot 10^{4}$. This shows that the presence of an external bump, induced by environmental effect, would switch on potentially large tidal effects in the inspiral part of the signal.
@@
However, in the final inspiral phases, the environment around each PBH may be disrupted because of tidal interactions. This disruption typically occurs when the binary’s semi-major axis is comparable to the Roche radius $R_{\text{\tiny Roche}}$ of the system Shapiro and Teukolsky (1983), which is defined as the distance at which the matter shell, held together solely by its gravitational force $F_{g}$ to the BH, will break apart due to the tidal force $F_{t}$ of a companion. This happens when the corresponding forces, acting on a volume element of mass $u$, are equal to each other, that is
@@
When the two masses are equal, one naturally obtains $R_{\text{\tiny Roche}}\simeq L$. In this estimate, we have considered an idealized scenario where the secondary BH’s environment is negligible compared to the secondary BH itself, because of its low density, destroying the ring surrounding its companion much before the frequency corresponding to the innermost stable circular orbit $f_{\text{\tiny ISCO}}\simeq 4.4\,{\rm kHz}\,M_{\odot}/(m_{1}+m_{2})$ is reached.
@@ III Observational prospects
We estimate the observational prospects of measuring tidal effects for subsolar binaries at current and future GW detectors using the Fisher Information Matrix (FIM) approach, and discuss the relevance of taking into account potential environmental effects when performing matched-filter search in the data to avoid a reduction in sensitivity of the detectors.
@@ III.1 Simulations and methods
Overall, the waveform model we adopt depends on 8 parameters $\vec{\theta}=\{{\cal M},\eta,\chi_{s},\chi_{a},t_{c},\phi_{c},\tilde{\Lambda},f_{\text{\tiny r}}\}$, where $(t_{c},\phi_{c})$ are the coalescence time and phase, set to $t_{c}=\phi_{c}=0$, and we have focused on optimally oriented binaries to eliminate the four angular parameters. The results of the Fisher forecast for these hyperparameters allow eventually to get information also on the forecasted uncertainties on both the individual BH masses $(m_{1},m_{2})$ and on the bump model parameters $(\epsilon,\tilde{L})$, as we will show in Sec. III.2.
@@ III.3 Detectability
Finally, let us comment on the prospects of detectability of subsolar PBH binaries when evolving in the vacuum or in an external environment. The faithfulness analysis reveals the degree to which a waveform model differs from another, such as comparing the presence or absence of the thin matter shell around each PBHs. High faithfulness indicates that the models closely match each other across a broad frequency range, while low faithfulness may suggest discrepancies in the waveform templates, possibly leading to systematic and wrong interpretations of the considered GW event.
@@ IV Conclusions
We conclude with a few words of caution. First of all, we should stress that our modeling of the disruption of the environment is somehow simplified. As the matter shell is disturbed, energy is taken from the binary system. Therefore, one expects the binary evolution to be impacted through e.g. dynamical friction. Additionally, as the mass surrounding the object is dispersed, one may also expect signatures of the environmental effects to be captured by an effective time-dependent binary mass. While these phenomena are present in principle, we stress that the mass of the matter shell we assume here is suppressed by the combination $\epsilon\tilde{L}$ compared to the PBH one. We leave an estimation of these additional phenomena for future work. Eventually, the signal of a “dirty” PBH binary should be distinguished from the one of neutron stars, or other exotic compact objects, characterized by an intrinsic TLN. The different morphology characterizing the GWs from the two systems provides us with the opportunity to achieve this. In the first case, tidal effects are present in the early phase of inspiral, while leaving a “naked” PBH binary closer to the ISCO frequency. In the latter case, the tidal effects are always present, until disruption of the compact object themselves Crescimbeni et al. (2024).

### v2

# Flea on the elephant: Tidal Love numbers in subsolar primordial black hole searches

@@ I Introduction
The detection of a subsolar mass object in a black hole (BH) merger is regarded as one of the definitive proofs of the primordial origin of a binary Green and Kavanagh (2021). However, leveraging this opportunity requires the ability to differentiate such an event from other astrophysical systems and possible new physics candidates Cardoso and Pani (2019); Barsanti et al. (2022); Franciolini et al. (2022); Crescimbeni et al. (2024).
@@
A key distinction between material compact objects and BHs is that the tidal deformability – measured by tidal Love numbers (TLNs) – of the latter is (supposed to be) zero in four-dimensional space-times Deruelle and Piran (1984); Binnington and Poisson (2009); Damour and Nagar (2009); Damour and Lecian (2009); Pani et al. (2015a, b); Gürlebeck (2015); Porto (2016); Le Tiec and Casals (2021); Chia (2021); Le Tiec et al. (2021); Hui et al. (2021); Charalambous et al. (2021a, b); Creci et al. (2021); Bonelli et al. (2022); Ivanov and Zhou (2023a); Charalambous et al. (2022); Katagiri et al. (2023a); Ivanov and Zhou (2023b); Berens et al. (2023); Bhatt et al. (2023); Sharma et al. (2024); Rai and Santoni (2024). This unique property of BHs in general relativity can be explained by special symmetries of the equations of motion for static and linear perturbations Hui et al. (2021); Charalambous et al. (2021a, b); Hui et al. (2022a, b); Charalambous et al. (2022); Ivanov and Zhou (2023b); Katagiri et al. (2023a); Bonelli et al. (2022); Kehagias et al. (2023); Ben Achour et al. (2022); Berens et al. (2023); Rai and Santoni (2024); Kehagias et al. (2024); Sharma et al. (2024) (see also Refs. De Luca et al. (2023a); Riva et al. (2024); Ivanov et al. (2024); Nair et al. (2023); Saketh et al. (2024); Perry and Rodriguez (2023); Chakraborty et al. (2024); De Luca et al. (2024) for recent studies discussing the role of nonlinearities and time-dependent tidal fields), but it does not apply to any other material object Porto (2016); Cardoso et al. (2017).
@@
This property is however invalidated in scenarios involving the presence of a cosmological constant Nair et al. (2024), in theories of modified gravity Cardoso et al. (2017, 2018); De Luca et al. (2023b); Barura et al. (2024) or higher dimensions Kol and Smolkin (2012); Cardoso et al. (2019); Hui et al. (2021); Rodriguez et al. (2023); Charalambous and Ivanov (2023); Charalambous (2024a, b); Ma et al. (2024), and especially in the presence of an external environment. The latter significantly alters the nature of the tidal interactions, potentially leading to nonzero TLNs for nonvacuum binary BHs. Examples of this effect include BHs surrounded by clouds of ultralight bosonic fields generated through accretion or superradiance Baumann et al. (2019); De Luca and Pani (2021); De Luca et al. (2023c); Brito and Shah (2023), or by fluids of matter and accretion disks Cardoso and Duque (2020); Cardoso et al. (2022); Cannizzaro et al. (2024).
@@
Since TLNs affect the gravitational wave (GW) signal of binary systems at high post-Newtonian (PN) order Flanagan and Hinderer (2008); Hinderer et al. (2010) and could be measured with good accuracy in the presence of large tidal interactions Chia et al. (2024a, b), they provide a natural key feature to test the nature of the compact object.
@@
Suppose now that in the future a subsolar merger event is detected with a large TLN. One would be immediately driven to think that such an event does not involve a primordial black hole (PBH). We show in this paper that this is not the case. Indeed, even a small environmental disturbance (the “flea”) to a BH (the “elephant”) generates a large and measurable effective TLN of the system. Such a “flea on the elephant” effect, here named in analogy with the quantum mechanical effect Graffi et al. (1984), is not surprising, it appears as well in the spectrum of the quasinormal modes of the GWs emitted during the ringdown phase following the merger of two BHs when tiny perturbations are present Jaramillo et al. (2021); Destounis et al. (2021); Cheung et al. (2022); Gasperin and Jaramillo (2022); Boyanov et al. (2023); Jaramillo (2022); Kyutoku et al. (2023); Sarkar et al. (2023); Destounis et al. (2024); Areán et al. (2023); Cownden et al. (2024); Destounis and Duque (2023); Courty et al. (2023); Boyanov et al. (2024); Cao et al. (2024); Cardoso et al. (2024); Ianniccari et al. (2024).
@@ II An Effective tidal Love number
The Dirac delta potential $V_{1}$, describing the external bump, is located at a distance $L$ much larger than $r_{s}$, and is the same considered in the aforementioned literature regarding the quasinormal modes Jaramillo et al. (2021); Destounis et al. (2021); Cheung et al. (2022); Gasperin and Jaramillo (2022); Boyanov et al. (2023); Jaramillo (2022); Kyutoku et al. (2023); Sarkar et al. (2023); Destounis et al. (2024); Areán et al. (2023); Cownden et al. (2024); Destounis and Duque (2023); Courty et al. (2023); Boyanov et al. (2024); Cao et al. (2024); Cardoso et al. (2024); Ianniccari et al. (2024), sharing strong similarities with the two rectangular barriers model of Refs. Chandrasekhar and Detweiler (1975); Barausse et al. (2014); Cheung et al. (2022). It is supposed to mimic the presence of a matter shell around the BH, with density profile characterised by a Poschl-Teller-like peak Cheung et al. (2022). The density amplitude is ${\cal O}(\epsilon)$, with width ${\cal O}(r_{s})\ll L$ and corresponding mass $\sim\epsilon\left(L/r_{s}\right)m$, where $m$ is the BH mass. This simple toy model allows to have an admittedly partial, but good, description of environmental effects present around the BH system Baumann et al. (2019); De Luca and Pani (2021); De Luca et al. (2023c); Brito and Shah (2023); Cardoso and Duque (2020); Cardoso et al. (2022); Cannizzaro et al. (2024), and has provided valuable insights on their role for BH spectroscopy, at least from an analytical perspective. It mimics for example a layer of a more complicated matter distribution which might spread over larger distances. The total effect of such matter distribution could thus be described as an integrated effect over many shells of this form. Our results are therefore intended to be conservative as far as the TLN is concerned, since tidal interactions could be effective for a longer portion of the inspiral (see discussion of tidal disruption below). The source creating the tidal force is assumed to be at distances $\gg L$, see Fig. 1 for a pictorial illustration.
@@
Thus, the TLNs of PBHs surrounded by weak rings could be sufficiently large and potentially impact the evolution of coalescing binaries. To get the feeling of the numbers, for the characteristic values $\epsilon\sim 10^{-4}$ and $L\sim 30\,r_{s}$, one obtains sizeable TLNs as $k_{2}\sim-1.5\cdot 10^{4}$. This shows that the presence of an external bump, induced by environmental effect, would switch on potentially large tidal effects in the inspiral part of the signal. It is worth commenting that a similar computation has been performed in Ref. Katagiri et al. (2023b) for a Gaussian small bump. The size and dependence of the corresponding TLNs on the bump parameters are compatible with ours, when overlap is possible.
@@
However, in the final inspiral phases, the environment around each PBH may be disrupted because of tidal interactions. This disruption typically occurs when the binary’s semi-major axis is comparable to the Roche radius $R_{\text{\tiny Roche}}$ of the system Shapiro and Teukolsky (1983), which is defined as the distance at which the matter shell, held together solely by its gravitational force $F_{g}$ to the BH, will break apart due to the tidal force $F_{t}$ of a companion. This happens when the corresponding forces, acting on a volume element of mass $u$, are equal to each other, which is
@@
When the two masses are equal, one naturally obtains $R_{\text{\tiny Roche}}\simeq L$. In this estimate, we have considered an idealized scenario where the secondary BH’s environment is negligible compared to the secondary BH itself, because of its low density, destroying the ring surrounding its companion much before the frequency corresponding to the innermost stable circular orbit (ISCO), $f_{\text{\tiny ISCO}}\simeq 4.4\,{\rm kHz}\,M_{\odot}/(m_{1}+m_{2})$, is reached.
@@ III Observational prospects
We estimate the observational prospects of measuring tidal effects for subsolar binaries at current and future GW detectors using the Fisher information matrix (FIM) approach, and discuss the relevance of taking into account potential environmental effects when performing a matched-filter search in the data to avoid a reduction in sensitivity of the detectors.
@@ III.1 Simulations and methods
Overall, the waveform model we adopt depends on eight parameters $\vec{\theta}=\{{\cal M},\eta,\chi_{s},\chi_{a},t_{c},\phi_{c},\tilde{\Lambda},f_{\text{\tiny r}}\}$, where $(t_{c},\phi_{c})$ are the coalescence time and phase, set to $t_{c}=\phi_{c}=0$, and we have focused on optimally oriented binaries to eliminate the four angular parameters. The results of the Fisher forecast for these hyperparameters allow eventually to get information also on the forecasted uncertainties on both the individual BH masses $(m_{1},m_{2})$ and on the bump model parameters $(\epsilon,\tilde{L})$, as we will show in Sec. III.2.
@@ III.3 Detectability
Finally, let us comment on the prospects of detectability of subsolar PBH binaries when evolving in the vacuum or in an external environment. The faithfulness analysis reveals the degree to which a waveform model differs from another, such as comparing the presence or absence of the thin matter shell around each PBH. High faithfulness indicates that the models closely match each other across a broad frequency range, while low faithfulness may suggest discrepancies in the waveform templates, possibly leading to systematic and wrong interpretations of the considered GW event.
@@ IV Conclusions
We conclude with a few words of caution. First of all, we should stress that our modeling of the disruption of the environment is somehow simplified. As the matter shell is disturbed, energy is taken from the binary system. Therefore, one expects the binary evolution to be impacted through e.g. dynamical friction. Additionally, as the mass surrounding the object is dispersed, one may also expect signatures of the environmental effects to be captured by an effective time-dependent binary mass. While these phenomena are present in principle, we stress that the mass of the matter shell we assume here is suppressed by the combination $\epsilon\tilde{L}$ compared to the PBH one. We leave an estimation of these additional phenomena for future work. Eventually, the signal of a “dirty” PBH binary should be distinguished from the one of neutron stars, or other exotic compact objects, characterized by an intrinsic TLN. The different morphology characterizing the GWs from the two systems provides us with the opportunity to achieve this. In the first case, tidal effects are present in the early phase of inspiral, while leaving a “naked” PBH binary closer to the ISCO frequency. In the latter case, the tidal effects are always present, until disruption of the compact objects themselves Crescimbeni et al. (2024).